If the probability density of X is given by f(x) = = { 2.-3 for 2 > 1 for r < 1, check whether its mean and variance exist.
Answer #1

The given function is - f(x) = 2x for xx1 for a 21 O Now, Mean = E(X) x. 2x3 dhe 」 > 2x2 dhe ག - 2 - Again, Variance = E(x-4)